OpenStudy (awesome781):

Scenario 1: Imagine that an economy produces two goods, flashlights and fishing lures. In 2011, the economy produced 70 flashlights and 40 fishing lures, and the prices of flashlights and fishing lures were $5 and $12, respectively. In 2012, the economy produced 85 flashlights and 50 fishing lures, and the prices of flashlights and fishing lures were $7 and $15, respectively. Based on the information in Scenario 1, nominal GDP grew by about ________ percent from 2011 to 2012. 23 31 62 162

OpenStudy (awesome781):

@mary.rojas

OpenStudy (mary.rojas):

@nincompoop

OpenStudy (awesome781):

@Abhisar

OpenStudy (abhisar):

I am really sorry, but this is not my forte, my help in this can prove disastrous for uh \(\color{green}{\huge\ddot\smile}\)

OpenStudy (mary.rojas):

am sorry :( hopefully someone else can help you

OpenStudy (awesome781):

i know someon who can help me but the problem is he is offline

OpenStudy (awesome781):

@annas

OpenStudy (ankit042):

Initial sum = 70*5+12*40= 830 next year = 85*7 + 50*15= 1345 Change = 1345-830/830

OpenStudy (ankit042):

@Awesome781 can you solve now?

OpenStudy (awesome781):

I got 1344

OpenStudy (ankit042):

No it is (1345-830)/830

OpenStudy (ankit042):

(1345-830) is the change and then you divide it with previous year to get the %change

OpenStudy (awesome781):

ok now I got a decimal answer O.O

OpenStudy (ankit042):

Yeah and now you have to round off. Check the question it says "grew by about" where about means to round off to nearest whole number

OpenStudy (awesome781):

oh lol I knew that :P I got C. 62

OpenStudy (ankit042):

Good!

OpenStudy (awesome781):

Thanks for the help @ankit042 you are the besttt!!!!!

OpenStudy (ankit042):

haha you are welcome!